OpenStudy (anonymous):

For a function f(x)=b^x, if 0

OpenStudy (anonymous):

Lots of things you can say here. One that comes to mind is "approach zero as x grows larger."

OpenStudy (campbell_st):

well the range will be 1< f(x) < b for the given domain

OpenStudy (anonymous):

The answer choices are A. Rise from right to left B.rise from left to right C.fall from right to left D.fall from left to right

OpenStudy (anonymous):

Without giving away the answer, I suppose I could give a hint: the thing I said is another way of saying one of those things.

OpenStudy (anonymous):

I choose c is that correct ?

OpenStudy (anonymous):

No, you have it backwards. As x increases (moving left to right on the graph), f(x) decreases to 0 (falls).

OpenStudy (campbell_st):

the behaviour of the curve is dependent upon the value of b is b > 1 then it rises from left to right e.g. y = 2^x is b = 1 the curve is simply the line y = 1 if b < 1 then the curve falls from left to right... e.g.y = (0.5)^x